Find the slope of a line parallel to the line that passes through the following points. (4,0) and (3.8,2)
Answer by nerdybill(7384)

Two lines are parallel if their slopes are the same.
Determine slope between:
(4,0) and (3.8,2)
.
m = (y2-y1)/(x2-x1)
m = (2-0)/(3.8-4)
m = (2)/(-0.2)
m = (1)/(-0.1)
m = (10)/(-1)
m = -10
.
answer: -10